OpenStudy (fabgirl1994):

Identify the vertex for the graph of y = −3x2 + 6x − 4. (1 point) (1, 5) (1, −1) (−1, −13) (−1, −9)

OpenStudy (s4sensitiveandshy):

you can find the x-coordinate of the vertex by using this formula -b/2a that will give you the x-coordinate. and then substitute x for its value into the equation for y-coordinate

OpenStudy (s4sensitiveandshy):

x-coordinate of the vertex represents the axis of symmetry

OpenStudy (scarlettfarra2000):

f(x)=−3x2+6x−2 Find the derivative of f(x) and set it to zero. Solve for x then. Afterwards if you want your y you simply plug the x you found to the original function f(x) The arrow points to the vertex. We can graph it and estimate where the vertex is but the best way to do this is by finding the derivative of the function....why does this work? Well the thing is that when you take the derivative we know that whatever value for x that makes the function zero will tell us where the curve changes. After we find the x all we can do is plug in the x we found to the original function

OpenStudy (scarlettfarra2000):

To find the vertex use the formula below to find the x vertex point of y = −3x^2 + 6x − 2 x=−b2a Plug our points in the the formula x=−(+62(−3) x=−6−6 x = 1 //This is our x vertex Now we have the x coord (1, ?) we need to find the y coord To find the y coord just plug our x coord, which is -1 into the equation y=−3x2+6x−2 y=−3(1)2+6(1)−2 y=−3+6−2 y=3−2 y =1 //This is our y vertex point Now we have x and y so our vertex is at (1, 1)

OpenStudy (anonymous):

(1,-1) Refer to the attachment.